The in-line fans for the K-hut and Strip and Wash room are now in place and working. This means that we can begin the process of commissioning the strip and Wash room. At the moment we still need a few more safety indicators on if the room has negative or positive pressure and to get the hot and cold water nozzles in place. This work seems to be moving forward at a good pace.
The replacement of the old truss sensors is moving forward. As mentioned before the old sensors have been removed and the new ones are being assembled.
Lift 3rd VSS platform into place, test fit, secured in place in prep for install.
Moved upper VSS platform into bay area, painted modified area and moved onto dome floor.
Work around the VIRUS Support Structure (VSS) continues. We have made some modifications to the work platforms and three of those have been test fit, secured in place and in preparation for the installation in two weeks. This week we did some work on possible coupling between the VSS and the structure. Our concern was that wind shake on the VSS would cause the mirrors to move. We did a number of test at the CCAS tower involving people physically moving the currently empty VSS enclosures. The concern seems to be warranted as the FWHM of the induced movement is about a quarter of an arc-second. We will look into identifying the members that cause the most coupling between the two and mitigate the motion.
The software team out here has been working on the DMI and Tip/Tilt to get it realigned with the new mount models that were installed last week.
The big news this week is the arrival of the HPF enclosure. Despite some very thick fog and misty rain we were able to get the truck unloaded and placed in the spectrograph room without tearing up the newly prepared floor. The PSU assembly team will arrive late next week.